Find the area of the triangle with vertices A(6,-7), B(11,-3) and
C(2,-2).

Answer by Alan3354(69443)

Find the area of the triangle with vertices A(6,-7), B(11,-3) and
C(2,-2).
----------
x: 6 11 +2 +6
y:-7 -3 -2 -7
----------------
Add the diagonal products starting at the upper left.
-18 - 22 - 14 = -54
---
Add the diagonal products starting at the lower left.
-77 - 6 - 12 = -95
-----
Area = 1/2 the difference = 41/2
Area = 20.5 sq units